# TechCoach: Towards Technical Keypoint-Aware Descriptive Action Coaching

### v1

@@ 3.3 Comparison with existing datasets
We compare our dataset with related AQA datasets. As shown in Tab. 1, the detailed keypoint-level commentary allows us to separate our EE4D-DAC dataset from the other datasets [22, 39, 15] with keypoint-level annotations on both type and scale. Furthermore, we conduct an evaluation to show the guiding ability of the instance-level commentary given by different datasets. Specifically, we prompt the ChatGPT to rate the commentary from 0 to 5 as the Coach-Score. As shown in the last column in Tab. 1, commentary MTL-AQA [27] and MTL-NAE [54] is not suit for coaching scenario as the original text mainly focuses on the action and score rather than action guiding and skill improvement.
For more details about the prompts, the pre-defined body parts, the examples and dataset comparison strategies, please refer to the Appendix.
@@ 4.1 Problem Formulation
Figure 3: An overview of the proposed TechCoach. Given an action video, TechCoach first extracts visual context embeddings with a pre-trained visual backbone and a Spatial-Temporal Context Enhancer. After that, a Context-aware Keypoint Reasoner is adopted to learn keypoint-related quality embeddings by querying visual context under the supervision of keypoint-level coaching commentary. Finally, prompted by the visual context and the keypoint-related quality embeddings, a unified Keypoint-aware Action Assessor is then employed to provide the overall coaching commentary together with the quality score. K.P.: Keypoints. Best viewed in color.
Figure 4: An overview of the Context-aware Keypoint Reasoner. Best viewed in color.
@@ 4.2 Overview
An overview of the proposed Technical Keypoint-aware Descriptive Action Coach (TechCoach) is shown in Fig. 3. The key idea of TechCoach is to mimic the reasoning process of human coaches by incorporating keypoint-level reasoning into the action coaching pipeline. Our framework consists of three parts. Firstly, given an input video $v$, a Visual Encoding Module will extract the visual context embedding. After that, a Context-aware Key-point Reasoner is adopted to take the visual context embedding and general technical keypoints as input to perceive the action quality on each keypoints and provide the key-point quality embeddings. Finally, a Unified Keypoint-aware Action Assessor is utilized to collaborate the visual context and keypoint-aware quality embeddings, then predict the final quality score and generate coaching commentary.
Discussions. Note that our model is based on the assumption that the general technical keypoints are available. We want to argue that this assumption is reasonable because: 1) It is aligned with real-world coaching scenarios that a coach is well-informed about the keypoints and able to reason through the keypoints and learner’s executions. 2) Obtaining keypoints is not resource-intensive as we can prompt LLMs (e.g., ChatGPT) to provide the keypoints.
@@ 4.3 Visual Encoding
4.4 Context-aware Key-point Reasoner
@@
To generate comprehensive coaching commentary for an action execution, an intuitive approach is to directly input the video (and the keypoint) features into the text generator. However, without explicit modeling, such an approach is unable to effectively perceive the relationship between visual context and the keypoints. To address this, TechCoach adopts a Context-aware Key-point Reasoner to learn keypoint-related quality embeddings by querying the video context embeddings under the supervision of the coaching commentary for each technical keypoint. An overview of Context-aware Keypoint Reasoner is shown in Fig. 4.
@@
- Keypoint Queries Construction. Without loss of generality, given the N general technical keypoints corresponding to the input video, we first extract the keypoint features $f_{kp}\in R^{N\times D}$ with a pre-trained text encoder (e.g., DeBERTa [11]) and a linear mapper. After that, we augment $f_{kp}$ by introducing two types of learnable prompts. Specifically, we use a group of keypoint-type prompts $f_{kt}\in R^{N\times D}$ to identify which dimension (e.g., head & eyes, human-object interaction) each keypoint feature belongs to. Besides, we adopt two quality-related prompts $f_{q}\in R^{2\times D}$ to indicate the aspects of strength and weakness. Based on the definitions of the learnable prompts, the augmented keypoint features $f^{*}_{kp}\in R^{N\times 2\times D}$ are computed as:
$f^{*}_{kp}[i,j]=f_{kp}[i]+f_{kt}[i]+f_{q}[j].\vspace{-0.2cm}$ (1)
@@
Such an augmentation allows us to inject the keypoint and quality information into the keypoint queries.
@@
- Keypoint-driven Quality Reasoning. Our goal here is to decouple the visual context features corresponding to the keypoints and quality aspects. To this end, we regard the augmented keypoint features $f^{*}_{kp}$ as the queries and visual context features $f_{v}$ as the keys and values, then utilize layers of multi-head cross attention module [36] to obtain the keypoint-related quality embeddings $f_{kq}\in R^{N\times 2\times D}$. $f_{kq}[i,j]$ indicates the action quality on keypoint $i$ and aspect $j$ ($1$ for the strength and $2$ for the weakness).
@@
To ensure $f_{kq}$ carries the quality information related to the action execution on the technical keypoints, our idea is to align the keypoint-related quality embeddings $f_{kq}$ with the features of the corresponding keypoint-level coaching commentary. To this end, a keypoint-level alignment loss is adopted, which can be written as:
$L_{Align}=\sum^{N}_{i=1}\sum^{2}_{j=1}w_{kc}[i,j]*D(\psi(f_{kq}[i,j]),f_{kc}[i,j]),\vspace{-0.2cm}$ (2)
@@
where $f_{kc}$ indicates the features of keypoint-level commentary extracted by the text encoder; $\psi$ is a linear mapper that projects the $f_{kq}$ back to the dimension of the text features; $w_{kc}[i,j]$ is a binary scalar to filter those cases where commentary on some aspects is not provided (see Discussions in Sec. 3.2); $D(\cdot,\cdot)$ indicates a distance measurement between two features, and we use the L2 Distance by default.
4.5 Unified Keypoint-aware Action Assessor
@@
After obtaining the visual context embeddings and the keypoint-related quality embeddings, we employ a Transformer-based Unified Keypoint-aware Action Assessor (KA2) to jointly predict the action quality score and generate the overall coaching commentary.
@@
Specifically, the input sequence to the KA2 module is constructed by concatenating the following components: (1) The generated text tokens $T_{1:i-1}$; (2) a learnable score-prediction token $f_{s}$; (3) the flattened $f_{kc}$ as keypoint tokens; (4) the flattened $f_{v}$ as video tokens. For generating commentary, we follow [17, 54] by employing a Mask-Token Prediction supervision ($L_{MTM}$) during training and using the Next-Token Prediction paradigm during inference. For score regression, the score-prediction token from the output layer is passed to a fully connected regressor, optimized using a Mean Squared Error (MSE) loss ($L_{MSE}$).
@@
Moreover, to guide the KA2 module in progressively integrating information in a sequence from low-level to high-level (i.e., starting with visual context level, then keypoint level, and finally the decision level), a progressive action coaching attention mask is further adopted on the KA2, which is shown in the lower right part of Fig. 3. Specifically, each keypoint token independently attends to itself and the video tokens. Besides, the score-prediction token attends to itself and all the keypoint and video tokens. For commentary generation, KA2 integrates all information by attending to the generated text tokens, the score-prediction token, all the keypoint tokens, and the video tokens.
@@
$L=L_{MTM}+\lambda_{1}L_{MSE}+\lambda_{2}L_{Align},\vspace{-0.2cm}$ (3)
@@
Methods Years Score Regression
$\rho\uparrow$ RL2$\downarrow$
@@
InternVideo2-MLP - 69.75 4.70
USDL [35] CVPR21 70.81 4.51
SwinBERT [17] CVPR22 71.37 4.59
PGMI [54] CVPR24 70.93 4.38
TechCoach(Ours) 72.46 4.35
@@
CoRe [50] w/ 10 Exem. ICCV21 73.17 4.35
TPT[3] w/ 10 Exem. ECCV22 73.77 4.29
(a) Comparison with existing methods on score regression.
Methods Years Commentary Generation
@@
General Multimodal Large Language Models
VideoLLaVA-7B [16] ACL24 22.94 1.79 9.80 53.76 16.62 14.82
VideoChat2-7B [14] CVPR24 23.07 3.20 11.94 60.79 17.82 25.99
InternVideo2-Stage3-8B [41] ECCV24 27.52 2.98 12.83 63.80 37.85 28.86
InternVL2-8B [6] 2024 12.51 0.10 16.81 62.30 47.76 34.01
Task-Specific Models
SwinBERT[17] CVPR22 36.11 11.70 16.01 66.03 46.56 34.57
PGMI [54] CVPR24 36.78 14.00 16.24 66.42 49.02 37.04
TechCoach(Ours) - 37.06 14.62 16.39 66.89 50.44 38.15
(b) Comparison with existing methods on coaching commentary generation.
Table 2: Comparison with existing models. Our TechCoach outperforms both the general MLLMs and Multi-Task Learning-based AQA models on coaching commentary generation. Besides, TechCoach shows strong score regression abilities with the highest performance among the methods using single instance for predicting the score. “w/ 10 Exem.”: We gray out models that infer with 10 extra exemplars for fair comparisons.
@@ 5.1 Dataset and Evaluation Metrics
- Dataset. We conduct experiments on EE4D-DAC to provide a new benchmark and evaluate the performance of the proposed TechCoach. Following the official split of EgoExo4D [10], we separate our EE4D-DAC into training and evaluation set (including 3769 and 1074 instances, respectively).
- Evaluation Metrics. For score prediction, we use the Spearman’s rank correlation coefficient ($\rho$) and Relative-L2 Distance (RL2) as the metrics. For natural language generation, we adopt NLP metrics, including BERT [56], BLEU(B)[25], METEOR(M)[4] and CIDEr(C) [37] scores. Moreover, for better comparison with open-source MLLMs, we design two more LLM-based Metrics: (1) a Mention Score(GPT-M) to evaluate whether the generated commentary mentions the same technical details as in the ground truth; (2) a Quality Score(GPT-Q) to evaluate whether the generated commentary shares the same praises or improvement opinions as in the ground truth on those both-mentioned technical details. For more details about the evaluation metrics, please refer to the Appendix.
@@ 5.2 Implementation Details
For each video instance, we sample frames and segment them into 32 8-frame clips. Subsequently, we use pretrained InternVideo2 [41] to extract clip features and adopt average pooling to reduce the spatial and temporal resolutions. obtaining video features with a size of $16(T)\times 8(H)\times 8(W)$. We use the pretrained DeBERTa-xlarge [11] as the text encoder. We use both the ego- and exo-centric (best-exo in EgoExo4D) videos as input. We concatenate the multi-view visual features on channel dimension and feed them into a linear projector before going through the Spatial-Temporal Context Enhancer. For more implementation details, please refer to the Appendix.
@@ 5.3 Main Results
We compare the proposed methods with two branches of baselines: (1) Task-Specific Models: We select AQA- and VideoCaption-specific baselines including USDL [35], CoRe [50], TPT[3], SwinBERT[17] and PGMI [54]. For fair comparisons, we evaluate all domain-specific models by using the same visual backbone (i.e., InternVideo2) to extract features and the same fusion strategy on multi-view features. (2) General Multi-modal Large Language Models: We also select popular open-source MLLMs as the baselines, including VideoLLaVA [16], VideoChat2 [14], InternVideo2-Stage3 [41], and InternVL2 [6]. For more details about the baseline models, please refer to the Appendix. The comparison results are shown in Tab. 2 and we have the following observations:
Variants Score Commentary
@@
70.78 4.45 - - - -
- - 36.71 13.70 16.39 66.77
70.23 4.77 36.42 14.26 16.09 65.72
72.46 4.35 37.06 14.62 16.39 66.89
Table 3: Ablation studies on training losses. The results show that keypoint-level alignment is a core design to achieve a better trade-off between score regression and coaching commentary generation.
Variants Score Commentary
@@
w/o KP-Align 70.23 4.77 36.42 14.26 16.09 65.72
KP-CLS 72.13 4.45 36.78 13.69 16.19 64.69
KP-Align 72.46 4.35 37.06 14.62 16.39 66.89
Table 4: Ablation studies on the alternative solutions of keypoint-level alignment. The results show that the proposed keypoint-level alignment (KP-Align) cannot be simply replaced by the naive classification-based solution (KP-CLS).
1) Score Regression: Compared with Direct Regression-based (i.e., InternVideo2-MLP and USDL) and Multi-Task Learning-based (i.e., SwinBERT, PGMI) AQA methods, TechCoach achieves the best performance on score regression. Note that CoRe and TPT adopt a Contrastive Regression framework that utilizes 10 extra exemplars during inference, which brings stronger score regression ability.
2) Coaching Commentary Generation: Although recently proposed Multimodal Large Language Models (MLLMs) show great visual context understanding and reasoning abilities, most of them [41, 14, 16] still struggle to handle realistic coaching scenarios (under-perform the task-specific models). Notably, the best among the evaluated MLLMs is the InternVL2 [6], which shows comparable performance with the task-specific models on LLM-based metrics. While InternVL2 is able to mention similar technical details as the ground truth (with a 47.76 GPT-M score), it falls short in providing precise guiding opinions (only 34.01 on the GPT-Q score). Among all the compared methods, our TechCoach achieves the best performance among all the metrics on coaching commentary generation.
Variants Score Commentary
@@
70.83 4.49 36.25 13.34 15.88 66.36
71.96 4.47 35.99 13.24 15.95 66.51
72.00 4.39 36.43 12.57 16.07 66.72
72.46 4.35 37.06 14.62 16.39 66.89

### v2

@@ 3.3 Comparison with existing datasets
We compare our dataset with related AQA-related datasets. As shown in Sec. 3.1, the detailed TechPoint-level commentary allows us to distinguish our EE4D-DescCoach dataset from the other related datasets [25, 42, 16]. Furthermore, we conduct an evaluation to show the guiding ability of the instance-level commentary given by different datasets. Specifically, we prompt the ChatGPT to rate the commentary from 0 to 5 as the Coach-Score. As shown in the last column in Sec. 3.1, commentary MTL-AQA [30] and MTL-NAE [59] is not suit for coaching scenario as the original text mainly focuses on the action and score rather than action guiding and skill improvement. For more details about the prompts, the pre-defined body parts, the examples and quality ensurance, please refer to the Appendix.
@@ 4.1 Problem Formulation
Figure 3: An overview of the proposed TechCoach. Given an action video, TechCoach first extracts visual context embeddings with a pre-trained visual backbone and a Spatial-Temporal Context Enhancer. After that, a Context-aware Tech-point Reasoner is adopted to learn Tech-point-related quality embeddings by querying visual context under the supervision of TechPoint-level coaching commentary. Finally, prompted by the visual context and the TechPoint-related quality embeddings, a unified TechPoint-aware Action Assessor is then employed to provide the overall coaching commentary together with the quality score. T.P.: TechPoints. Best viewed in color.
Figure 4: An overview of the Context-aware TechPoint Reasoner. Best viewed in color.
@@ 4.2 Overview
An overview of the proposed Technical-Point-aware Descriptive Action Coach (TechCoach) is shown in Fig. 3. The key idea of TechCoach is to mimic the reasoning process of human coaches by incorporating TechPoint-level reasoning into the action coaching pipeline. Our framework consists of three parts. Firstly, given an input video $v$, a Visual Encoding Module will extract the visual context embedding. After that, a Context-aware TechPoint Reasoner is adopted to take the visual context embedding and general technical points as input to perceive the action quality on each TechPoints and provide the TechPoint quality embeddings. Finally, a Unified TechPoint-aware Action Assessor is utilized to collaborate the visual context and TechPoint-aware quality embeddings, then predict the final quality score and generate coaching commentary.
Discussions. Note that our model is based on the assumption that the general TechPoints are available. We want to argue that this assumption is reasonable because: 1) It is aligned with real-world coaching scenarios that a coach is well-informed about the TechPoints and able to reason through the TechPoints and learner’s executions. 2) Obtaining TechPoints is not resource-intensive as we can prompt LLMs (e.g., ChatGPT) to provide the TechPoints.
@@ 4.3 Visual Encoding
4.4 Context-aware TechPoint Reasoner
@@
To generate comprehensive coaching commentary for an action execution, an intuitive approach is to directly input the video (and the TechPoint) features into the text generator. However, without explicit modeling, such an approach is unable to effectively perceive the relationship between visual context and the TechPoints. To address this, TechCoach adopts a Context-aware TechPoint Reasoner to learn TechPoint-related quality embeddings by querying the video context embeddings under the supervision of the coaching commentary for each TechPoint. An overview of Context-aware TechPoint Reasoner is shown in Fig. 4.
@@
- TechPoint Queries Construction. Without loss of generality, given the N general TechPoints corresponding to the input video, we first extract the TechPoint features $f_{tp}\in R^{N\times D}$ with a pre-trained text encoder (e.g., DeBERTa [11]) and a linear mapper. After that, we augment $f_{tp}$ by introducing two types of learnable prompts. Specifically, we use a group of TechPoint-type prompts $f_{tt}\in R^{N\times D}$ to identify which dimension (e.g., head & eyes, human-object interaction) each TechPoint feature belongs to. Besides, we adopt two quality-related prompts $f_{q}\in R^{2\times D}$ to indicate the aspects of strength and weakness. Based on the definitions of the learnable prompts, the augmented TechPoint features $f^{*}_{tp}\in R^{N\times 2\times D}$ are computed as:
$f^{*}_{tp}[i,j]=f_{tp}[i]+f_{tt}[i]+f_{q}[j].\vspace{-0.2cm}$ (1)
@@
Such an augmentation allows us to inject the TechPoint and quality information into the TechPoint queries.
@@
- TechPoint-driven Quality Reasoning. Our goal here is to decouple the visual context features corresponding to the TechPoints and quality aspects. To this end, we regard the augmented TechPoint features $f^{*}_{tp}$ as the queries and visual context features $f_{v}$ as the keys and values, then utilize layers of multi-head cross attention module [39] to obtain the TechPoint-related quality embeddings $f_{tq}\in R^{N\times 2\times D}$. $f_{tq}[i,j]$ indicates the action quality on TechPoint $i$ and aspect $j$ ($1$ for the strength and $2$ for the weakness).
@@
To ensure $f_{tq}$ carries the quality information related to the action execution on the TechPoints, our idea is to align the TechPoint-related quality embeddings $f_{tq}$ with the features of the corresponding TechPoint-level coaching commentary. To this end, a TechPoint-level alignment loss is adopted, which can be written as:
$L_{Align}=\sum^{N}_{i=1}\sum^{2}_{j=1}w_{tc}[i,j]*D(\psi(f_{tq}[i,j]),f_{tc}[i,j]),\vspace{-0.4cm}$ (2)
@@
where $f_{tc}$ indicates the features of TechPoint-level commentary extracted by the text encoder; $\psi$ is a linear mapper that projects the $f_{tq}$ back to the dimension of the text features; $w_{tc}[i,j]$ is a binary scalar to filter those cases where commentary on some aspects is not provided (see Discussions in Sec. 3.2); $D(\cdot,\cdot)$ indicates a distance measurement between two features and we use the L2 Distance by default.
4.5 Unified TechPoint-aware Action Assessor
@@
After obtaining the visual context embeddings and the TechPoint-related quality embeddings, we employ a Transformer-based Unified TechPoint-aware Action Assessor (TA2) to jointly predict the action quality score and generate the overall coaching commentary.
@@
Specifically, the input sequence to the TA2 module is constructed by concatenating the following components: (1) The generated text tokens $T_{1:i-1}$; (2) a learnable score-prediction token $f_{s}$; (3) the flattened $f_{tq}$ as TechPoint tokens; (4) the flattened $f_{v}$ as video tokens. For generating commentary, we follow [19, 59] by employing a Mask-Token Prediction supervision ($L_{MTM}$) during training and using the Next-Token Prediction paradigm during inference. For score regression, the score-prediction token from the output layer is passed to a fully connected regressor, optimized using a Mean Squared Error (MSE) loss ($L_{MSE}$).
@@
Moreover, to guide the TA2 module in progressively integrating information in a sequence from lower to higher level (i.e., starting with visual context level, then TechPoint level, and finally the decision level), a progressive action coaching attention mask is further adopted on the TA2, which is shown in the lower right part of Fig. 3. Specifically, each TechPoint token independently attends to itself and the video tokens. Besides, the score-prediction token attends to itself and all the TechPoint tokens and video tokens. For commentary generation, TA2 integrates all information by attending to the generated text tokens, the score-prediction token, all the TechPoint tokens, and the video tokens.
@@
$L=L_{MTM}+\lambda_{1}L_{MSE}+\lambda_{2}L_{Align},\vspace{-0.3cm}$ (3)
@@
{NiceTabular}
l|c@  |c@   c@    Methods Visual Score Regression
Backbone $\rho\uparrow$ RL2$\downarrow$
@@
InternVideo2-MLP IV2 [44] 69.75 4.70
USDL [38] IV2 [44] 70.81 4.51
SwinBERT [19] IV2 [44] 71.37 4.59
PGMI [59] IV2 [44] 70.93 4.38
TechCoach(Ours) IV2 [44] 72.46 4.35
@@
CoRe [54] w/ 10 Exem. IV2 [44] 73.17 4.35
TPT[3] w/ 10 Exem. IV2 [44] 73.77 4.29
@@
(b) Comparison with existing methods on coaching commentary generation. Table 2: Comparison with existing models. Our TechCoach outperforms both the general MLLMs and Multi-Task Learning-based AQA models on coaching commentary generation. Besides, TechCoach shows strong score regression abilities with the highest performance among the methods using single instance for predicting the score. “w/ 10 Exem.”: We gray out models that infer with 10 extra exemplars for fair comparisons. “Txt.Gen.”: Text Generator. “IN2”: InternVideo2.
@@ 5.1 Dataset and Evaluation Metrics
- Dataset. To build new benchmark and evaluate the effectiveness of our TechCoach, we separate EE4D-DescCoach into training and evaluation set (with 3769 and 1074 instances) following the official split of EgoExo4D [10]. - Evaluation Metrics. For score prediction, we use the Spearman’s rank correlation coefficient ($\rho$) and Relative-L2 Distance (RL2) as the metrics. For natural language generation, we adopt NLP metrics, including BERT [61], BLEU(B)[28], METEOR(M)[4] and CIDEr(C) [40] scores. Moreover, for better comparison with open-source MLLMs, we design two more LLM-based Metrics: (1) a Mention Score(GPT-M) to evaluate whether the generated commentary mentions the same technical details as in the ground truth; (2) a Quality Score(GPT-Q) to evaluate whether the generated commentary shares the same praises or improvement opinions as in the ground truth on those both-mentioned technical details (i.e., action quality perception ability). Please refer to Appendix for more details.
@@ 5.2 Implementation Details
For each video instance, we sample frames and segment them into 32 8-frame clips. Subsequently, we use pretrained InternVideo2 [44] to extract clip features and adopt average pooling to reduce the spatial and temporal resolutions. obtaining video features with a size of $16(T)\times 8(H)\times 8(W)$. We use both the ego- and exo-centric (best-exo in EgoExo4D) videos as input. We concatenate the multi-view visual features on channel dimension and feed them into a linear projector before going through the Spatial-Temporal Context Enhancer. We use similar multi-modal Transformer encoder as in [59, 19] to integrate inputs and generate text. More implementation details are provided in the Appendix.
@@ 5.3 Main Results
As shown in Sec. 4.5, we compare the proposed methods with two branches of baselines: (1) Task-Specific Models: We select task-specific baselines including USDL [38], CoRe [54], TPT[3], SwinBERT[19] and PGMI [59]. For fair comparisons, we evaluate all task-specific models by using the same visual backbone (i.e., InternVideo2) to extract features and the same fusion strategy on multi-view features. (2) General Multi-modal Large Language Models: We compare popular open-source MLLMs including VideoLLaVA [18], VideoChat2 [15], InternVideo2-S3 [44], and InternVL2 [6] under zero-shot evaluation settings. We provide more details about the baselines in the Appendix.
{NiceTabular} ccc|cc|cccc Variants Score Commentary
@@
✓ 70.78 4.45 - - - -
✓ - - 36.71 13.70 16.39 66.77
✓ ✓ 70.23 4.7736.42 14.26 16.09 65.72
✓ ✓ ✓ 72.46 4.35 37.06 14.62 16.39 66.89
Table 3: Ablation studies on training losses. TechPoint-level alignment is a core design to achieve a better trade-off between score regression and coaching commentary generation.
Refer to caption Figure 5: A further illustration of the impact of TechPoint-level alignment loss. After adding $L_{Align}$, a much stronger improvement is observed on action quality perception ability (evaluated by GPT-Q Score). Please refer to Sec. 5.4 for more analysis.
- Score Regression: Compared with Direct Regression-based (i.e., InternVideo2-MLP and USDL) and Multi-Task Learning-based (i.e., SwinBERT, PGMI) AQA methods, TechCoach achieves the best performance on score regression. Note that CoRe and TPT adopt a Contrastive Regression framework that utilizes 10 extra exemplars during inference, thereby enhancing their score regression capability.
@@
- More Comparison Results. In Appendix, we also provide deeper explorations and analysis on various aspects, including (i) User study, (ii) Adding general TechPoints into the prompts, (iii) Abnormal performance of InternVL2 on traditional language metrics, (iv) Fine-tuned MLLM.
{NiceTabular} l|cc|cccc Variants Score   Commentary
@@
w/o TP-Align 70.23 4.77 36.42 14.26 16.09 65.72
TP-CLS 72.13 4.45 36.78 13.69 16.19 64.69
TP-Align 72.46 4.35 37.06 14.62 16.39 66.89
Table 4: Ablation studies on the alternative solutions of TechPoint-level alignment. The results show that the proposed TechPoint-level alignment (TP-Align) cannot be simply replaced by the naive classification-based solution (TP-CLS).
{NiceTabular} ccc|cc|cccc Variants Score Commentary
[…]
✓ 70.83 4.49 36.25 13.34 15.88 66.36
✓ 71.96 4.47 35.99 13.24 15.95 66.51
✓ ✓ 72.00 4.39 36.43 12.57 16.07 66.72
✓ ✓ ✓ 72.46 4.35 37.06 14.62 16.39 66.89